OpenStudy (anonymous):

Find the surface-area-to-volume ratio of a right circular cylinder. (Hint: For a right circular cylinder, S

OpenStudy (anonymous):

\[2 \pi r^2+ 2 \pi r h\]= surface area

OpenStudy (anonymous):

\[\pi r^2h\]= volume

OpenStudy (anonymous):

\[\frac{ 2 \pi r^2+ 2 \pi r h }{ \pi r^2 h }= \frac{ 2 \pi r(r+h) }{ \pi r^2 h }=\frac{2( r+h )}{ rh}\]